Suppose you draw a card, put it back in the deck, and draw another one. What is the probability that the first card is a three and the second is a five? (Enter your probability as a fraction.)
Answer by greenestamps(13198)



The probability of a 3 on the first draw is 4/52 = 1/13; since the first card is put back in the deck, the probability of a 5 on the second draw is 4/52 = 1/13.

The probability of both is the product of the two probabilities: 1/13*1/13 = 1/169.